# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ 4 Two-sample testing
In this section, we focus on the non-parametric two-sample problem. In Section 4.1, we formally define the two-sample testing framework, explain how permutations or a wild bootstrap can be used to simulate the null, and present non-asymptotic level guarantees. In Section 4.2 and Section 4.3, we present power guarantees in terms of MMD and $L^{2}$ Sobolev uniform separation rates, respectively, covering standard, efficent, differentially private and robust testing frameworks. We refer the reader to Schrab (2025b, Section 2.1) for a detailed introduction to the Maximum Mean Discrepancy (MMD).
@@ 4.1 Framework, bootstrap and level
 [^6] A more general framework is the one of credal two-sample testing, see Chau et al. (2025) for details. Given independent i.i.d. samples $X_{1},\dots,X_{m}$ from a distribution $P$, and i.i.d. samples $Y_{1},\dots,Y_{n}$ from a distribution $Q$, the aim is to test whether the two distributions are equal, that is, ${\mathcal{H}}_{0}\colon P=Q$, or not, i.e., ${\mathcal{H}}_{1}\colon P\neq Q$. Following the general hypothesis testing notation of Section 1, this corresponds to having ${\mathcal{P}}$ as the space of all pairs of distributions, ${\mathcal{P}}_{0}$ as $\{(P,Q)\in{\mathcal{P}}:P=Q\}$, and ${\mathcal{P}}_{1}$ as $\{(P,Q)\in{\mathcal{P}}:P\neq Q\}$. We use the notation ${\mathbb{X}}_{m}\coloneqq(X_{1},\dots,X_{m})$, ${\mathbb{Y}}_{n}\coloneqq(Y_{1},\dots,Y_{n})$ and let $N=\min(m,n)$.
@@
Two-sample testing can be framed as testing for exchangeability (Section 1). Permuted two samples are constructed by, first, combining the two original samples, permuting all the elements, and then splitting them again into two separate samples (of the original sample sizes). Samples which come from the same distribution (i.e., null hypothesis) are exactly the ones which are exchangeable. Indeed, it is clear that null samples, which are i.i.d., are exchangeable. Under the alternative, both permuted samples can be seen as drawn i.i.d. from a mixture of the two distinct distributions, this shows that the original and permuted samples do not have the same distribution (i.e., samples are not exchangeable). Hence, two-sample testing corresponds to testing for exchangeability, and the two-sample null hypothesis can be simulated using permutations.
@@
For notation purposes, let $Z_{i}=X_{i}$ for $i=1,\dots,n$ and $Z_{i}=Y_{i-m}$ for $i=n+1,\dots,m+n$. As aforementioned, given a permutation $\pi$ of $\{1,\dots,m+n\}$, permuting the original samples $\mathbb{X}_{m}$ and $\mathbb{Y}_{n}$ with respect to $\pi$ leads to the permuted samples $\mathbb{X}_{m}^{\pi}\coloneqq(Z_{\pi(1)},\dots,Z_{\pi(m)})$ and $\mathbb{Y}_{n}^{\pi}\coloneqq(Z_{\pi(m+1)},\dots,Z_{\pi(m+n)})$. Given any statistic function $T$, the test statistic is simply $T(\mathbb{X}_{m},\mathbb{Y}_{n})$, and permuted statistics can be computed as $T(\mathbb{X}_{m}^{\pi},\mathbb{Y}_{n}^{\pi})$ for various permutations $\pi$ randomly sampled. A test can then be constructed using these permutations and can be performed efficiently, as explained in Section 1.2 and Schrab et al. (2023, Appendix C). This resulting test is well-calibrated with non-asymptotic level $\alpha$ as desired (Romano and Wolf, 2005a, Lemma 1, see also Kim and Schrab, 2023, Lemma 15).
@@
While the permutation approach presented above is applicable when using any type of MMD estimators (and more generally when using any other statistic). We now consider the wild bootstrap method presented in details in Section 1.2, which is specifically designed for (MMD) estimators expressed as one-sample second-order statistics (see Schrab, 2025b, Section 3). The wild bootstrapped statistics are computed as ${|{\mathcal{D}}|}^{-1}\sum_{(i,j)\in{\mathcal{D}}}\varepsilon_{i}\varepsilon_{j}h(X_{i},X_{j})$ where $\varepsilon_{1},\dots,\varepsilon_{n}$ are realisations of i.i.d. Rademacher variables. Using these allows to construct a test following the procedure of Section 1.2, the resulting test is guaranteed to control the type I error asymptotically (Chwialkowski et al., 2014, Theorem 1). In the two-sample setting with $m=n$, the MMD estimator (e.g., Schrab, 2025b, Equation 11) admits as core function $h^{\operatorname{MMD}}_{k}(x,x^{\prime};y,y^{\prime})\coloneqq k(x,x^{\prime})-k(x^{\prime},y)-k(x,y^{\prime})+k(y,y^{\prime})$, which satisfies $h^{\operatorname{MMD}}_{k}(x,y^{\prime};x^{\prime},y)=-h^{\operatorname{MMD}}_{k}(x,x^{\prime};y,y^{\prime})$. Leveraging this identity allows to prove that, when $m=n$, using a wild bootstrap is equivalent to using permutations which are only able to swap $X_{i}$ and $Y_{i}$ (or not) for $i=1,\dots,n$ (Schrab et al., 2023, Appendix B). This guarantees non-asymptotic type I error control of the wild bootstrap MMD test, which can be implemented efficiently (Schrab et al., 2023, Appendix C). See Section 1.2 for computational details, as well as for a discussion on when to use each of the two bootstrapping methods.
@@
The permutation-based MMD, dpMMD and dcMMD tests (Section 3), as well as the wild bootstrap MMD test, all tightly control the probability of type I error by $\alpha$ at every sample size as desired (Schrab et al., 2023, Proposition 1; Kim and Schrab, 2023, Theorem 5; Schrab and Kim, 2025, Lemmas 1 and 4). This non-asympotic level is preserved when using efficient estimators (Schrab et al., 2022b, Proposition 1), as well as when using adaptivity over kernels, either via pooling (properties of the permutation method, Romano and Wolf, 2005a, Lemma 1, combined with Schrab et al., 2023, Proposition 1; see also discussion around Biggs et al., 2023, Theorem 1) or via aggregation (Schrab et al., 2023, Proposition 8).
@@
The MMD-based tests depend on the choice of kernel, which in practice greatly impacts the test power. To solve this problem, kernel adaptivity can be performed either via aggregation (Section 2) or via kernel pooling (Schrab, 2025b, Section 4). In practice, we recommend using the MMDAgg test (Schrab et al., 2023) and the normalised MMDFuse test (Biggs et al., 2023).
@@ 4.2 Uniform power against alternatives separated in MMD metric
For a fixed kernel $k$, the MMD test is powerful provided that (Section A.1 and Kim and Schrab, 2023, Theorem 7)
$\mathrm{MMD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}}$ (13)
which is minimax optimal (Kim and Schrab, 2023, Theorem 8). When using unnormalised kernel pooling (Schrab, 2025b, Section 4) with the mean pooling function, by leveraging the linearity of the discrepancy in the kernel (Schrab, 2025b, Equation 75), we obtain the uniform separation rate (Section A.4)
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ \mathrm{MMD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}}.$ (14)
The MMD test with unnormalised kernel pooling (Schrab, 2025b, Section 4) with kernel pooling function either max or fuse (with fusing parameter $\nu\geq\max(N,\log(|{\mathcal{K}}|))$), achieves the uniform separation rate (Section A.4 and Biggs et al., 2023, Theorems 2 and 3)
$\max_{k\in{\mathcal{K}}}\,\mathrm{MMD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big{\}}}{N}}$ (15)
where a typical choice is $|{\mathcal{K}}|=\log(N)$ (e.g., Schrab et al., 2023, Corollary 10) leading to the rate $(N/\log\log N)^{-1/2}$ with an iterated logarithmic cost for adaptivity. We note that for the fuse case, it is possible to get the fuse pooling function on the left hand side of Equation 15 instead of the maximum, but using the relation of Schrab (2025b, Equation 78) between both quantities, we can replace this fuse by the maximum provided that the fusing parameter $\nu$ is greater than both $N$ and $\log(|{\mathcal{K}}|)$, which is almost always the case in practice. In the remaining of this subsection, we present results only for fuse and max kernel pooling (variants of Equation 15) without always mentioning the weak assumption $\nu\geq\max(N,\log(|{\mathcal{K}}|))$, while results for kernel pooling with the mean function still hold with the exact same rate similarly to Equation 14 but are not explicitly presented.
@@
The test using a block MMD B-statistic (Schrab, 2025b, Equation 70), consisting of $B$ blocks, controls the type II error by $\beta$ if (Section A.5)
$\mathrm{MMD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{B\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}}$ (16)
when using a fixed kernel $k$, and if (Section A.4)
$\max_{k\in{\mathcal{K}}}\,\mathrm{MMD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{B\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big{\}}}{N}}$ (17)
when using a pooled unnormalised kernel collection ${\mathcal{K}}$ (Schrab, 2025b, Section 4) with max or fuse pooling functions. It is possible to choose ${\mathcal{K}}$ such that $|{\mathcal{K}}|=\log(|{\mathcal{D}}|/N)\approx\log(N/B)$ (Schrab et al., 2022b, Theorem 2.ii) since the associated design ${\mathcal{D}}$ is of size $B\lfloor N/B\rfloor^{2}\asymp N^{2}/B$. When using a complete statistic (i.e., with $B=1$ block), the uniform separation rate achieved is the minimax one as in the standard testing setting above. As $B$ increases from $1$ to $N$, Equation 16 quantifies how the uniform separation rate deteriorates from being minimax to finally not even converging to zero. This highlights the trade-off between computational efficiency and power (speed of uniform separation rate).
@@
The $(\varepsilon,\delta)$-differentially private dpMMD test (Kim and Schrab, 2023, Algorithm 1 and Section 4) achieves the uniform separation rate (Section A.2 and Kim and Schrab, 2023, Theorem 7)
$\mathrm{MMD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\Biggl{\{}\sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}},\,\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N\xi}\Biggr{\}}$ (18)
@@
The dcMMD (Schrab and Kim, 2025, Algorithm 1 and Section 3) and dpMMD (Schrab and Kim, 2025, Algorithm 2, Section 5, Appendix E) tests with fixed kernel, designed to be robust against corruption of up to $r$ samples, are guaranteed to be powerful as soon as (Section A.3 and Schrab and Kim, 2025, Theorems 1.i and 3)
$\mathrm{MMD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\biggl{\{}\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta)\}}{N}},\ \frac{r}{N}\biggr{\}}$ (19)
@@
$\max_{k\in{\mathcal{K}}}\,\mathrm{MMD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\biggl{\{}\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta),\log(|{\mathcal{K}}|)\}}{N}},\ \frac{r}{N}\biggr{\}}.$ (20)

### v2

@@ 4 Two-sample testing
In this section, we focus on the non-parametric two-sample problem. In Section˜4.1, we formally define the two-sample testing framework, explain how permutations or a wild bootstrap can be used to simulate the null, and present non-asymptotic level guarantees. In Section˜4.2 and Section˜4.3, we present power guarantees in terms of MMD and $L^{2}$ Sobolev uniform separation rates, respectively, covering standard, efficent, differentially private and robust testing frameworks. We refer the reader to Schrab (2025b, Section 2.1) for a detailed introduction to the Maximum Mean Discrepancy (MMD).
@@ 4.1 Framework, bootstrap and level
 [^6] A more general framework is the one of credal two-sample testing, see Chau et al. (2025) for details.Given independent i.i.d. samples $X_{1},\dots,X_{m}$ from a distribution $P$, and i.i.d. samples $Y_{1},\dots,Y_{n}$ from a distribution $Q$, the aim is to test whether the two distributions are equal, that is, ${\mathcal{H}}_{0}\colon P=Q$, or not, i.e., ${\mathcal{H}}_{1}\colon P\neq Q$. Following the general hypothesis testing notation of Section˜1, this corresponds to having $\mathcal{P}$} as the space of all pairs of distributions, ${\mathcal{P}}_{0}$ as $\{(P,Q)\in{\mathcal{P}}:P=Q\}$, and ${\mathcal{P}}_{1}$ as $\{(P,Q)\in{\mathcal{P}}:P\neq Q\}$. We use the notation ${\mathbb{X}}_{m}\coloneqq(X_{1},\dots,X_{m})$, ${\mathbb{Y}}_{n}\coloneqq(Y_{1},\dots,Y_{n})$ and let $N=\min(m,n)$.
@@
Two-sample testing can be framed as testing for exchangeability (Section˜1). Permuted two samples are constructed by, first, combining the two original samples, permuting all the elements, and then splitting them again into two separate samples (of the original sample sizes). Samples which come from the same distribution (i.e., null hypothesis) are exactly the ones which are exchangeable. Indeed, it is clear that null samples, which are i.i.d., are exchangeable. Under the alternative, both permuted samples can be seen as drawn i.i.d. from a mixture of the two distinct distributions, this shows that the original and permuted samples do not have the same distribution (i.e., samples are not exchangeable). Hence, two-sample testing corresponds to testing for exchangeability, and the two-sample null hypothesis can be simulated using permutations.
@@
For notation purposes, let $Z_{i}=X_{i}$ for $i=1,\dots,n$ and $Z_{i}=Y_{i-m}$ for $i=n+1,\dots,m+n$. As aforementioned, given a permutation $\pi$ of $\{1,\dots,m+n\}$, permuting the original samples $\mathbb{X}_{m}$ and $\mathbb{Y}_{n}$ with respect to $\pi$ leads to the permuted samples $\mathbb{X}_{m}^{\pi}\coloneqq(Z_{\pi(1)},\dots,Z_{\pi(m)})$ and $\mathbb{Y}_{n}^{\pi}\coloneqq(Z_{\pi(m+1)},\dots,Z_{\pi(m+n)})$. Given any statistic function $T$, the test statistic is simply $T(\mathbb{X}_{m},\mathbb{Y}_{n})$, and permuted statistics can be computed as $T(\mathbb{X}_{m}^{\pi},\mathbb{Y}_{n}^{\pi})$ for various permutations $\pi$ randomly sampled. A test can then be constructed using these permutations and can be performed efficiently, as explained in Section˜1.2 and Schrab et al. (2023, Appendix C). This resulting test is well-calibrated with non-asymptotic level $\alpha$ as desired (Romano and Wolf, 2005a, Lemma 1, see also Kim and Schrab, 2023, Lemma 15).
@@
While the permutation approach presented above is applicable when using any type of MMD estimators (and more generally when using any other statistic). We now consider the wild bootstrap method presented in details in Section˜1.2, which is specifically designed for (MMD) estimators expressed as one-sample second-order statistics (see Schrab, 2025b, Section 3). The wild bootstrapped statistics are computed as ${|{\mathcal{D}}|}^{-1}\sum_{(i,j)\in{\mathcal{D}}}\varepsilon_{i}\varepsilon_{j}h(X_{i},X_{j})$ where $\varepsilon_{1},\dots,\varepsilon_{n}$ are realisations of i.i.d. Rademacher variables. Using these allows to construct a test following the procedure of Section˜1.2, the resulting test is guaranteed to control the type I error asymptotically (Chwialkowski et al., 2014, Theorem 1). In the two-sample setting with $m=n$, the MMD estimator (e.g., Schrab, 2025b, Equation 11) admits as core function $h^{\operatorname{MMD}}_{k}(x,x^{\prime};y,y^{\prime})\coloneqq k(x,x^{\prime})-k(x^{\prime},y)-k(x,y^{\prime})+k(y,y^{\prime})$, which satisfies $h^{\operatorname{MMD}}_{k}(x,y^{\prime};x^{\prime},y)=-h^{\operatorname{MMD}}_{k}(x,x^{\prime};y,y^{\prime})$. Leveraging this identity allows to prove that, when $m=n$, using a wild bootstrap is equivalent to using permutations which are only able to swap $X_{i}$ and $Y_{i}$ (or not) for $i=1,\dots,n$ (Schrab et al., 2023, Appendix B). This guarantees non-asymptotic type I error control of the wild bootstrap MMD test, which can be implemented efficiently (Schrab et al., 2023, Appendix C). See Section˜1.2 for computational details, as well as for a discussion on when to use each of the two bootstrapping methods.
@@
The permutation-based MMD, dpMMD and dcMMD tests (Section˜3), as well as the wild bootstrap MMD test, all tightly control the probability of type I error by $\alpha$ at every sample size as desired (Schrab et al., 2023, Proposition 1; Kim and Schrab, 2023, Theorem 5; Schrab and Kim, 2025, Lemmas 1 and 4). This non-asympotic level is preserved when using efficient estimators (Schrab et al., 2022b, Proposition 1), as well as when using adaptivity over kernels, either via pooling (properties of the permutation method, Romano and Wolf, 2005a, Lemma 1, combined with Schrab et al., 2023, Proposition 1; see also discussion around Biggs et al., 2023, Theorem 1) or via aggregation (Schrab et al., 2023, Proposition 8).
@@
The MMD-based tests depend on the choice of kernel, which in practice greatly impacts the test power. To solve this problem, kernel adaptivity can be performed either via aggregation (Section˜2) or via kernel pooling (Schrab, 2025b, Section 4). In practice, we recommend using the MMDAgg test (Schrab et al., 2023) and the normalised MMDFuse test (Biggs et al., 2023).
@@ 4.2 Uniform power against alternatives separated in MMD metric
For a fixed kernel $k$, the MMD test is powerful provided that (Section˜A.1 and Kim and Schrab, 2023, Theorem 7)
$\mathrm{MMD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}}$ (13)
which is minimax optimal (Kim and Schrab, 2023, Theorem 8). When using unnormalised kernel pooling (Schrab, 2025b, Section 4) with the mean pooling function, by leveraging the linearity of the discrepancy in the kernel (Schrab, 2025b, Equation 75), we obtain the uniform separation rate (Section˜A.4)
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ \mathrm{MMD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}}.$ (14)
The MMD test with unnormalised kernel pooling (Schrab, 2025b, Section 4) with kernel pooling function either max or fuse (with fusing parameter $\nu\geq\max(N,\log(|{\mathcal{K}}|))$), achieves the uniform separation rate (Section˜A.4 and Biggs et al., 2023, Theorems 2 and 3)
$\max_{k\in{\mathcal{K}}}\,\mathrm{MMD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big\}}{N}}$ (15)
where a typical choice is $|{\mathcal{K}}|=\log(N)$ (e.g., Schrab et al., 2023, Corollary 10) leading to the rate $(N/\log\log N)^{-1/2}$ with an iterated logarithmic cost for adaptivity. We note that for the fuse case, it is possible to get the fuse pooling function on the left hand side of Equation˜15 instead of the maximum, but using the relation of Schrab (2025b, Equation 78) between both quantities, we can replace this fuse by the maximum provided that the fusing parameter $\nu$ is greater than both $N$ and $\log(|{\mathcal{K}}|)$, which is almost always the case in practice. In the remaining of this subsection, we present results only for fuse and max kernel pooling (variants of Equation˜15) without always mentioning the weak assumption $\nu\geq\max(N,\log(|{\mathcal{K}}|))$, while results for kernel pooling with the mean function still hold with the exact same rate similarly to Equation˜14 but are not explicitly presented.
@@
The test using a block MMD B-statistic (Schrab, 2025b, Equation 70), consisting of $B$ blocks, controls the type II error by $\beta$ if (Section˜A.5)
$\mathrm{MMD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{B\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}}$ (16)
when using a fixed kernel $k$, and if (Section˜A.4)
$\max_{k\in{\mathcal{K}}}\,\mathrm{MMD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{B\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big\}}{N}}$ (17)
when using a pooled unnormalised kernel collection $\mathcal{K}$} (Schrab, 2025b, Section 4) with max or fuse pooling functions. It is possible to choose $\mathcal{K}$} such that $|{\mathcal{K}}|=\log(|{\mathcal{D}}|/N)\approx\log(N/B)$ (Schrab et al., 2022b, Theorem 2.ii) since the associated design $\mathcal{D}$} is of size $B\lfloor N/B\rfloor^{2}\asymp N^{2}/B$. When using a complete statistic (i.e., with $B=1$ block), the uniform separation rate achieved is the minimax one as in the standard testing setting above. As $B$ increases from $1$ to $N$, Equation˜16 quantifies how the uniform separation rate deteriorates from being minimax to finally not even converging to zero. This highlights the trade-off between computational efficiency and power (speed of uniform separation rate).
@@
The $(\varepsilon,\delta)$-differentially private dpMMD test (Kim and Schrab, 2023, Algorithm 1 and Section 4) achieves the uniform separation rate (Section˜A.2 and Kim and Schrab, 2023, Theorem 7)
$\mathrm{MMD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \max\Biggl\{\sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}},\,\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N\xi}\Biggr\}$ (18)
@@
The dcMMD (Schrab and Kim, 2025, Algorithm 1 and Section 3) and dpMMD (Schrab and Kim, 2025, Algorithm 2, Section 5, Appendix E) tests with fixed kernel, designed to be robust against corruption of up to $r$ samples, are guaranteed to be powerful as soon as (Section˜A.3 and Schrab and Kim, 2025, Theorems 1.i and 3)
$\mathrm{MMD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \max\biggl\{\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta)\}}{N}},\ \frac{r}{N}\biggr\}$ (19)
@@
$\max_{k\in{\mathcal{K}}}\,\mathrm{MMD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \max\biggl\{\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta),\log(|{\mathcal{K}}|)\}}{N}},\ \frac{r}{N}\biggr\}.$ (20)